1÷[1+1÷{1+1÷(1+1÷3)}]
1
Expert's answer
2013-05-16T09:03:19-0400
ok, let's begin from the bottom (or the right side of our expression):
1÷[1+1÷{1+1÷(1+1÷3)}]=1÷[1+1÷{1+1÷(1+1/3)}]=1÷[1+1÷{1+1÷(4/3)}]=
Now we proceed "to the left" opening the braces:
=1÷[1+1÷{1+3/4}]=1÷[1+1÷{7/4}]=1÷[1+4/7]=1÷[11/7]=7/11 or in decimals it's approximately 0.64